Chapter 25

You have two flat metal plates, each of area 1.32 m2, with which to construct a parallel-plate capacitor. (a) If the capacitance of the device is to be 1.13 F, what must be the separation between the plates? (b) Could this capacitor actually be constructed?
Answer #1

(a)

we know that

C = \varepsilono*A/d

we can write

d = \varepsilono*A/C

d = 8.85*10-12 *1.32 / 1.13

d = 10.33*10-12 m

d = 1.033*10-11 m

(b)

Size of an atom in the order of 1*10-10 m . Here is the separation of atoms is ten time smaller than the atom in the plate .Hence,it would not be truly separated . Soa suitable capacitor could be constructed.